OpenStudy (anonymous):

The ends of the 0.57-m bar remain in contact with their respective support surfaces. End B has a velocity of 0.56 m/s and an acceleration of 0.57 m/s2 in the directions shown. Determine the angular acceleration α (positive if counterclockwise, negative if clockwise) of the bar and the acceleration of end A (positive if up, negative if down).

OpenStudy (anonymous):

ignore the first post

OpenStudy (anonymous):

the problem is in the image

OpenStudy (souvik):



OpenStudy (souvik):

now\[\frac{ AA'*\cos 30 }{ \Delta t }=\frac{ BB'*\cos 45 }{ \Delta t }\] so..\[\frac{ \sqrt{3} }{ 2 }v _{AA'}=\frac{ 1 }{ \sqrt{2} }v _{BB'}\]